OpenStudy (anonymous):

struggling to find the answer... i know what the answer is but cant seem to get there...... (x^1/2 + y^1/2)^2

OpenStudy (wolfe8):

First of all show me what you did. First is squaring it.

hero (hero):

\(\left(x^{1/2} + y^{1/2}\right)^2 = \left(\sqrt{x} + \sqrt{y}\right)^2\) \( = \left(\sqrt{x} + \sqrt{y}\right)\left(\sqrt{x} + \sqrt{y}\right)\) \(= \sqrt{x}\left(\sqrt{x} + \sqrt{y}\right) + \sqrt{y}\left(\sqrt{x} + \sqrt{y}\right)\) \(= x + \sqrt{xy} + \sqrt{xy} + y\) \(= x + 2\sqrt{xy} + y\)